OpenStudy (anonymous):

How do I simplify this? -9sqrt(28a^2)*1/3sqrt(63a) It equals -126 Sqrt[a] Sqrt[a^2]

OpenStudy (jim766):

\[-9\sqrt{28a ^{2}}* \frac{ 1 }{ 3 }\sqrt{63a}\] right?

OpenStudy (anonymous):

./facepalm. Ya, that is the right question. XD

OpenStudy (jim766):

\[-9*2a\sqrt{7}* \frac{ 1 }{ 3 }*3\sqrt{7a}\]

OpenStudy (jim766):

\[-18a \sqrt{7}* \sqrt{7a}\]

OpenStudy (jim766):

\[-126 \sqrt{a}\]

OpenStudy (anonymous):

Awesome, thanks a ton:)